OpenStudy (anonymous):

In a car dealership's lot, 6/24 of the cars are silver. Answers: A:20% B:25% C:52% D:64%

OpenStudy (jojo4eva):

B. 25%

OpenStudy (jojo4eva):

6/24=1/4=25%

OpenStudy (jojo4eva):

1/4=0.25*100=25%

OpenStudy (anonymous):

Thank you!:)